40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 6353     Entry time: Mon Mar 5 06:11:08 2012
Author: kiwamu 
Type: Summary 
Category: LSC 
Subject: plans 

Plans:

  •  DRMI (PRMI) + one arm test before the LVC meeting
  •  Study of the funny sensing matrix and the RAM offset effects before the LVC meeting
  •  Glitch hunting

Action items:

  • MC beam pointing 
    • to make the PZT1 pitch relax
  • OSA setup
    •    a long BNC cable for monitoring the signal in the control room
  • Power budget on the AP table
    • in order to ensure the laser power on each photo diode
  •  POP22/110 sideband monitor
    • installation of an RF amp
    • building a diplexer
    • connect the signals to the demod boards 
  •  Calibration of the demod boards
    • calibrate the conversion loss of the mixers to calibrate all the LSC signals to watts / meter
  •  (1+G) correction for the glitch time series data
  • Simulation study for the RAM offset
    • How much offset do we get due to the RAM ? and how do the offsets screw up the sensing matrix ?
  •  A complete set of the MICH characterization
    •   DC power
    •   Sensing matrix
    •   Noise budget
    •   OSA
    •   Estimation of the RAM offset 
    •  Summarize the results in the wiki
  •  A complete set of the PRMI/DRMI characterization
    •  The same stuff as the MICH characterization
  •  DRMI + one arm test
    •   Monitor the evolution of the sensing matrix during the arm is brought to the resonance

   
 

ELOG V3.1.3-